Integration Procedures (Integration by Parts)


∫ θ cosθ dθ


1
Expert's answer
2021-04-29T16:58:39-0400
"\int\theta\cos \theta d\theta"

Integration by Parts


"\int u dv=uv-\int v du"

"u=\theta , du=d\theta"


"dv=\cos \theta d\theta, v=\int \cos \theta d\theta=\sin \theta"

"\int\theta\cos \theta d\theta=\theta\sin \theta-\int\sin \theta d\theta"

"=\theta\sin \theta+\cos \theta +C"


"\int\theta\cos \theta d\theta=\theta\sin \theta+\cos \theta +C"